# Statistical Principles for Platform Trials

### v2

@@ 6 Conclusion
In conclusion, maintaining stringent control of error rates within clinical trials is crucial for ensuring the reliability of study outcomes and the safety of approved treatments. As clinical trial designs become increasingly complex, particularly with the rise of platform trials, the statistical community and regulatory bodies must continuously adapt and refine their methodologies to address these challenges effectively
@@ 7.1 Terms
  • ERpF: Error rate per family
  • FAR: False Approval Rate
  • FDR: False Discovery Rate
  • FWER: Family wise error rate
@@ 7.2 Notation
  • $\mu_{SC}$ Shared control sample mean
  • $\hat{\mu}_{SC}$ Shared control sample mean estimate
  • $V^{\star}$

### v3

@@ 6 Conclusion
In conclusion, maintaining stringent control of error rates within clinical trials is crucial for ensuring the reliability of study outcomes and the safety of approved treatments. As clinical trial designs become increasingly complex, particularly with the rise of platform trials, the statistical community and regulatory bodies must continuously adapt and refine their methodologies to address these challenges effectively.
@@ 7.1 Terms
  • ALS: Amyotrophic lateral sclerosis (ALS)/Lou Gehrig’s disease; a fatal progressive neuromuscular disease
  • ERpF: Error Rate per Family; measures errors across studies
  • ERFw: Error Rate Family wise; error within each study
  • FAR: False Approval Rate; proportion of drug compounds that lack efficacy incorrectly approved
  • FDR: False Discovery Rate; expected proportion of rejections that are incorrect
  • FWER: Family Wise Error Rate; probability of type I error in a family
@@ 7.2 Notation
  • $\mu_{SC}$: Shared control sample mean
  • $\hat{\mu}_{SC}$: Shared control sample mean estimate
  • $\hat{\mu}_{SC}^{\star}$: Standardized shared control sample mean estimate
  • $X$: Indicator that type I error is committed as a result of $i$th study
  • $V$: The number of Type I errors committed in a family
  • $Z$: Indicator that $i$th confidence interval erroneously shows efficacy
  • $V^{\star}$: The number of treatments with their confidence intervals for efficacy not covering their true parameter values given $n$ confidence intervals
  • $SE$: Standard error
    • Subscripts denote standard error of respective term
  • $k$: Number of studies
  • $n$: Maximum number of treatment arms erroneously approved
  • $T^{\star}_{i}$: Statistic for comparing $k$ new compounds against shared control
    $T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$
  • $R$: the number of correct rejection of false nulls
  • $n_{SC}$: sample size for shared control
  • $n_{i}$: sample size for $i$th treatment
  • $\lambda_{i}$: Have that
    $\lambda_{i}=\left(1+\frac{n_{SC}}{n_{i}}\right)^{-\frac{1}{2}}$
  • $\rho_{ij}$: Correlation between sample statistics $T_{i}$ and $T_{j}$
    • Note: Can be written as a product of $\lambda_{i}$ and $\lambda_{j}$.